Hello! I am a bit confused with this problem. PLEASE HELP ME.
find three consecutive even integers such that their sum decreased by the third is 82.

Answer by ankor@dixie-net.com(22740)

find three consecutive even integers
x, (x+2, (x+4)
:
Just write what it says," such that their sum decreased by the third is 82.
x + (x+2) + (x+4) - (x+4) = 82
2x + 2 = 82
2x = 82 - 2
2x = 80
x = 40
;
Check: 40 + 42 + 44 - 44 = 82